### v1

# Maintaining Plasticity in Deep Continual Learning

Abstract
Modern deep-learning systems are specialized to problem settings in which training occurs once and then never again, as opposed to continual-learning settings in which training occurs continually. If deep-learning systems are applied in a continual learning setting, then it is well known that they may fail catastrophically to remember earlier examples. More fundamental, but less well known, is that they may also lose their ability to adapt to new data, a phenomenon called loss of plasticity. We show loss of plasticity using the MNIST and ImageNet datasets repurposed for continual learning as sequences of tasks. In ImageNet, binary classification performance dropped from 89% correct on an early task down to 77%, or to about the level of a linear network, on the 2000th task. Such loss of plasticity occurred with a wide range of deep network architectures, optimizers, and activation functions, and was not eased by batch normalization or dropout. In our experiments, loss of plasticity was correlated with the proliferation of dead units, with very large weights, and more generally with a loss of unit diversity. Loss of plasticity was substantially eased by $L^{2}$-regularization, particularly when combined with weight perturbation (Shrink and Perturb). We show that plasticity can be fully maintained by a new algorithm—called continual backpropagation—which is just like conventional backpropagation except that a small fraction of less-used units are reinitialized after each example. This continual injection of diversity appears to maintain plasticity indefinitely.  [^1] Code is available at https://github.com/shibhansh/loss-of-plasticity
keywords:
continual learning, deep learning, loss of plasticity, ImageNet, MNIST, backpropagation
[…]
[inst1]organization=Department of Computing Science, University of Alberta
[…]
[inst2]organization=CIFAR AI Chair, Alberta Machine Intelligence Institute
1 Loss of Plasticity
Modern deep-learning systems have become specialized to problem settings in which training occurs once on a large data set and then never again. Such train-once settings were used in all the early successes of deep learning in speech recognition and image classification. Later, when deep learning was adapted to reinforcement learning (e.g., Mnih et al., 2015), techniques such as replay buffers and batching were introduced to make it very nearly a train-once setting. Recent applications of deep learning such as GPT-3 (Brown et al., 2020), and DallE (Ramesh et al., 2021) were also trained on data as a single large batch. Of course, in many applications the data distribution changes over time and training must continue in some form; the most common strategy in these cases has been to continually collect data and then, on occasion, train a new network from scratch, again in a train-once setting. The train-once setting has been integral to the design and practice of modern deep-learning methods.
In contrast, the continual learning problem setting focuses on learning continually from new data. The continual learning setting is advantageous for problems where the learning system faces a changing data stream. For example, consider a robot tasked with navigating a house. Under the train-once setting, the robot would have to be retrained from scratch or risk becoming obsolete whenever the house’s layout changed. If the layout changed frequently, then constant retraining from scratch would be required. Under the continual learning setting, on the other hand, the robot could simply learn from the new data and keep adapting to the changes in the house. Continual learning has drawn increasing attention in recent years, and new specialized meetings are being organized to focus on it, such as the Conference on Life-long Learning Agents (CoLLAS). In this paper, we focus on the continual learning setting.
Deep learning systems, when exposed to new data, tend to forget most of what they have previously learned, a phenomenon known as “catastrophic forgetting.” In other words, deep learning methods do not maintain stability in continual learning problems. This phenomenon was first shown in early neural networks in the late 1900s (McCloskey & Cohen, 1989; French, 1999). Recently, with the advent of deep learning, catastrophic forgetting has received renewed attention as many papers have been dedicated to maintaining stability in deep continual learning (Kirkpatrick et al., 2017; Yoon et al., 2018; Aljundi et al., 2019; Golkar et al., 2019; Rajasegaran et al., 2019; Javed & White, 2019).
Different from catastrophic forgetting, and even more fundamental to continual learning, is the ability to keep learning from new data. We refer to this ability as plasticity. Maintaining plasticity is essential for continual learning systems as it allows them to adapt to changes in their data streams. Continual learning systems that lose plasticity can become useless if their data stream changes. In this paper, we focus on the issue of loss of plasticity. This focus is different but complementary to the more-common focus on catastrophic forgetting.
Do deep-learning systems lose plasticity in continual learning problems? Some evidence that they do comes from the psychology literature of the early 2000s. Ellis & Lambon Ralph (2000), Zevin & Seidenberg (2002), and Bonin et al. (2004) showed loss of plasticity in early neural networks in supervised regression problems. These papers used a setup where a set of examples was presented to the network for a certain number of epochs, and then the training set was expanded with additional examples and training proceeded for another number of epochs. They found that the error for the examples in the original training set was lower than for the later-added examples after controlling for the number of epochs. These papers provided evidence that deep learning and the backpropagation algorithm (Rumelhart et al., 1986) on which it is founded tend to produce loss of plasticity. A limitation of these early works is that the networks used were relatively shallow by today’s standards, and the algorithms used are not those that are most popular today. For example, they did not use the ReLU activation function or the Adam optimizer. The early psychology research on artificial neural networks provides tantalizing hints, but does not provide a clear answer to the question of whether or not modern deep learning networks exhibit loss of plasticity.
Turning now to the machine learning literature, we can find several recent studies suggesting loss of plasticity in deep learning. Chaudhry et al. (2018) observed loss of plasticity in continual image classification problems. Their experiments are not fully satisfactory as a demonstration of loss of plasticity because they introduced other confounding variables. In their setup, when a new task was presented, new outputs, called heads, were added to the network; the number of outputs grew as additional tasks were encountered. Loss of plasticity was thereby confounded with the effects of interference from old heads. Chaudhry et al. found that when old heads were removed at the start of a new task, the loss of plasticity was minimal, which suggests that the loss of plasticity they observed was mainly due to interference from old heads. A second limitation of Chaudhry et al.’s study is that it used only ten tasks and thus did not assess loss of plasticity when deep learning methods face a long sequence of tasks.
Ash and Adams (2020) showed that initializing a neural network by first training it on a fraction of the dataset can lead to much lower final performance compared to the case in which the network is trained once on the entire dataset. The worse performance after pretraining is an important example of loss of plasticity in deep learning and hints at the possibility of a major weakness in full continual learning with many task changes or non-stationarity. Berariu et al. (2021) extended Ash and Adams’ work to show that the more pretraining stages there were, the worse the final performance. However, Berariu et al.’s experiments were performed in a stationary problem setting, and the loss of plasticity they observed was small. Although the results in these papers hint at a fundamental loss of plasticity in deep learning systems, none directly demonstrated loss of plasticity in continual learning.
More evidence for loss of plasticity in modern deep learning has been obtained in the reinforcement learning community, where recent papers have shown substantial loss of plasticity. Nishikin et al. (2022) showed that early learning in reinforcement learning problems could harm later learning, an effect they termed “primacy bias.” This result could be due to the loss of plasticity of deep learning networks in continual learning settings, as reinforcement learning is inherently continual due to changes in the policy. Lyle et al. (2022) also showed that some deep reinforcement learning agents could lose the ability to learn new functions over time. These are important data points, but the conclusions that can be drawn from them are limited by the inherent complexity of modern deep reinforcement learning. All of these papers, both from the psychology literature at the turn of the century and recent papers in machine learning and reinforcement learning, are evidence that deep-learning systems lose plasticity, but fall short of a fully satisfactory demonstration of the phenomena.
In this paper, we attempt a more definitive answer to the question of loss of plasticity in modern deep learning. We show that deep-learning methods suffer from loss of plasticity in continual supervised learning problems and that such a loss of plasticity can be severe. First, we demonstrate that deep learning suffers from loss of plasticity in a continual supervised-learning problem based on the ImageNet dataset that involves thousands of learning tasks. Using supervised learning tasks instead of reinforcement learning removes the complexity and associated confounding that inevitably arises in reinforcement learning. And having thousands of tasks allows us to assess the full extent of the loss of plasticity. Then we use two computationally cheaper problems (a variant of MNIST and the Slowly-Changing Regression problem) to establish the generality of deep learning’s loss of plasticity over a wide range of hyperparameters, optimizers, network sizes, and activation functions.
After showing the severity and generality of loss of plasticity in deep learning, we seek a deeper under-standing of its causes. This leads us to explore existing methods that might ameliorate loss of plasticity. We find that some methods, like Adam (Kingma & Ba, 2015) and dropout (Hinton et al., 2012), make the loss of plasticity significantly worse, while other methods, like $L^{2}$ regularization (Goodfellow et al., 2016, pp. 227-230) and Shrink and Perturb (Ash & Adams, 2020), substantially ease the loss of plasticity in many cases. Finally, we propose a new algorithm—continual backpropagation—that robustly solves the problem of loss of plasticity in both our systematic experiments with supervised learning problems and in a preliminary experiment with a reinforcement learning problem. The usual backpropagation algorithm has two crucial components: 1) initialization of the network’s connection weights with small random numbers and 2) stochastic gradient descent on every presentation of a training example (Rumelhart, Hinton & Williams, 1986). Continual backpropagation extends the initialization to all time steps, by selectively reinitializing a small fraction of low-utility hidden units at every presentation of a training example. Continual backpropagation performs both gradient descent and selective reinitialization continually.
2 Loss of Plasticity in ImageNet
Imagenet is a large database of images and their labels that has been influential throughout machine learning (Deng et al., 2009; Image-net.org) and that played a key role in the rise of deep learning (Krizhevsky et al., 2012). ImageNet allowed researchers to conclusively show that deep learning can solve a problem like object recognition, a hallmark of intelligence. Classical machine learning methods had a classification error rate of 25% (Russakovsky et al., 2015) when deep-learning methods reduced it to below 5% (Hu et al., 2018), the human error rate on the dataset.
The ImageNet database comprises millions of images labelled by nouns (classes) such as animals and everyday objects. In the standard train-once setting, the dataset is partitioned into a training set and a test set. The learning system is first trained on the training set, and then the learner is frozen and its performance is measured on the test set.
We sought to adapt Imagenet from the train-once setting to the continual-learning setting while mini-mizing all other changes. The Imagenet database we used consisted of 1000 classes, each of 700 images. We constructed from these a near-endless sequence of binary classification tasks. For example, the first task might be to distinguish Class1 from Class2, and the second task might be to distinguish Class3 from Class4. Each binary classification task was handled in conventional way. The 700 images for each class were divided into 600 images for a training set and 100 images for a test set. On each task, the deep learning network was first trained on the training set of 1200 images, and then its classification accuracy was assessed on the test set of 200 images. That training consisted of multiple passes through the training set, called epochs. After training and test on one task, the next task was begun based on a different pair of classes. By selecting the pairs of tasks randomly without replacement, we generated a sequence of 2000 different tasks. All tasks used the downsampled 32x32 version of the ImageNet dataset, as is often done to save computation (Chrabaszcz et al., 2017). We call the resulting continual-learning problem Continual ImageNet.
Network Architecture for Continual ImageNet
Layer 1: Convolutional + Max-Pooling
Number of Filters 32 Activation ReLU
[…]
Layer 2: Convolutional + Max-Pooling
Number of Filters 64 Activation ReLU
[…]
Layer 3: Convolutional + Max-Pooling
Number of Filters 128 Activation ReLU
[…]
Layer 4: Fully Connected
Output Size 128 Activation ReLU
Layer 5: Fully Connected
Output Size 128 Activation ReLU
Layer 6: Fully Connected
Output Size 2 Activation Linear
Table 1: Details of the artificial neural network used for the Continual ImageNet problem. The network has three convolutional layers followed by three fully connected layers.
We applied a wide variety of standard deep-learning networks to Continual ImageNet. We describe here results with a specific, representative network, using convolutional layers and the ReLU activation function. The network had three convolutional-plus-max-pooling layers followed by three fully connected layers, as detailed in Table 1. This network is narrower than those commonly used on the ImageNet dataset because here we are trying to discriminate only two classes at a time rather than all 1000. The final layer consisted of just two units, the heads, corresponding to the two classes. At task changes, the input weights of the heads were reset to zero. Resetting the heads in this way can be viewed as introducing new heads for the new classes. This resetting of the output weights is not ideal for studying plasticity, as the learning system gets access to privileged information on the timing of task changes (and we do not use it in the experiments reported later in this paper). We use it here because it is closest to the standard practice in deep continual learning, which is to introduce new heads for new classes.
We tested many variations of the network architecture and hyper-parameters to obtain good performance on the first binary classification task. The networks were trained using stochastic gradient descent (SGD) with momentum on the cross-entropy loss and initialized once prior to the first task. For each task, all learning algorithms performed 250 passes through the training set using mini-batches of size 100. The momentum hyperparameter was 0.9. We tested various step-size parameters but are only presenting the performance for step sizes 0.01, 0.001, and 0.0001 for clarity of the Figure. We performed 30 runs for each hyperparameter value, varying the sequence of tasks and other randomness. Across different hyperparameters, the same sequences of pairs of classes were used.
Figure 1: Loss of plasticity on a sequence of ImageNet binary classification tasks. The first plot shows performance over the first ten tasks, which sometimes improved initially before declining. The second plot shows performance over 2000 tasks, over which the loss of plasticity was extensive. The learning algorithm was backpropagation applied in the conventional deep-learning way.
As a measure of performance on a task, we used the percentage of the 200 test images that were correctly classified. This measure is plotted as a function of task number for the first ten tasks (left panel of Figure 1) and in bins of 50 tasks for all 2000 tasks (right panel of Figure 1) for the best step sizes. The first data point in the right panel is the average performance on the first 50 tasks; the next is the average performance over the next 50 tasks, and so on.
The largest step size, 0.01, performed the best on the first two tasks, but performance then fell on subsequent tasks, eventually reaching a level below the linear baseline. For even larger step sizes like 0.1, the network diverged after the first few tasks. At the smaller step sizes, performance rose initially and then fell and was only slightly better than the linear baseline after 2000 tasks. The performance for intermediate step-sizes like 0.003 followed the general trend of the ones presented in Figure 1. We have found this to be a common pattern in our experiments: for a well-tuned network, performance first improves, then falls substantially, ending near or below the linear baseline. We have observed this pattern for many network architectures, hyper-parameters, and optimizers. The specific choices of network architecture, hyper-parameters, and optimizers affect when the performance starts to drop, but we observed a severe performance drop for a wide range of choices.
This experiment was designed to closely follow standard deep-learning practice in every way possible, except in requiring the network to keep learning. Failure to learn better than a linear network in later tasks is substantial evidence that the standard practice of deep learning simply doesn’t work in continual problems. However, there are of course many other variations of deep learning that could be tried. To reach the next level of confidence and understanding we switch in the next section to a less computationally complex problem where we can study the phenomena more thoroughly.
3 Robust Loss of Plasticity in Permuted MNIST
MNIST (LeCun & Cortes, 1998) is one of the most common supervised-learning datasets used in deep learning. It consists of 60,000 28x28 grayscale images of the handwitten digits from 0 to 9 together with their correct labels. For example, the left image in Figure 2 shows an image that is labelled by the digit 7. The smaller number of classes and the simpler images enables much smaller deep learning networks to perform well on this dataset than are needed on ImageNet. The smaller networks in turn mean much less computation is needed to perform the experiments and thus many more and more varied experiments can be done, enabling us to perform deeper and more extensive studies of plasticity. As with ImageNet, MNIST is typically used in a train-once setting.
Figure 2: Left: An MNIST image with the label ‘7’; Right: A corresponding permuted image.
We created a continual supervised learning problem using permuted MNIST datasets (Goodfellow et al., 2014; Zenke et al., 2017). An individual permuted MNIST dataset is created by permuting the pixels in the original MNIST dataset. For example, the pixel that originally appeared in the upper-left corner might be moved to the lower-right corner, or to any other particular place in the image, and so on for all the other pixels. The right image in Figure 2 is an example of such a permuted image for an image labelled with the digit 7. Given a way of permuting, all 60,000 images are permuted in the same way to produce the new permuted MNIST dataset.
By repeatedly randomly selecting from the approximately $10^{1930}$ possible permutations, we created a sequence of 800 permuted MNIST datasets and supervised learning tasks. For each task, we presented each of its 60,000 images one-by-one in a random order to the learning network. Then we moved to the next permuted MNIST task and repeated the whole procedure, and so on for up to 800 tasks. No indication was given to the network at the time of task switching. With the pixels being permuted in a completely unrelated way, we might expect classification performance to fall dramatically at the time of each task switch. Nevertheless, across tasks, there could be some savings, some improvement of speed of learning, or alternatively there could be loss of plasticity—loss of the ability to learn across tasks. The network was trained on a single pass through the data and there are no mini-batches. We call this problem Online Permuted MNIST.
We applied feed-forward neural networks with three hidden layers to Online Permuted MNIST. We did not use convolutional layers, as they could not be helpful on the permuted problem (in MNIST, convolutional layers are often not used even on the standard, non-permuted problem). For each example, the network estimated the probabilities of each of the 10 classes (0–9), compared them to the correct label, and performed SGD on the cross-entropy loss. As a measure of performance, we recorded the percentage of times the network’s highest probability label was the correct over a tasks 60,000 images. We plot this per-task performance measure versus task number in the figures below. The weights were initialized according to a Kaiming distribution.
The first panel of Figure 3 shows the progression of performance across tasks for a network of with 2000 units per layer, and various values of the step-size parameter. Note that performance first increased across tasks, then began falling steadily across all subsequent tasks. These results are consistent with the loss of plasticity that we observed in ImageNet.
Figure 3: Loss of plasticity in Online Permuted MNIST is robust over step sizes, network sizes, and rates of change.
Next we varied the network size. Instead of 2000 units per layer, we tried 100, 1000, and 10,000 units per layer. In this experiment we ran for only 150 tasks, primarily because the largest network took much longer to run. The time courses of performance at good step sizes for each network size are shown in the middle panel of Figure 3. The loss of plasticity with continued training is most pronounced at the smaller network sizes, but even the largest networks showed some loss of plasticity.
Next we studied the effect of the rate at which the task changed. Going back to the original network with 2000-unit layers, instead of changing the permutation each 60,000 examples, we now changed it after each 10,000, 100,000, or 1M examples, and ran for 48M examples total no matter how often the task changed. The examples in these cases were selected randomly from the permuted MNIST task’s dataset, with replacement. As a performance measure of the network on a task, we used the percent correct over all of the task’s images. That is, there were only 48 data points at the slowest rate, and 4800 data points at the fastest rate. The progression of performance is shown in the right plot in Figure 3. Again performance fell across tasks, even if the change was very infrequent. Altogether these results show that the phenomena of loss of plasticity robustly arises in this form of backpropagation.
4 Loss of Plasticity With Different Activations in the Slowly Changing Regression Problem
There remains the issue of the network’s activation function. We have been using ReLU, the most popular choice at present, but there are a number of other possibilities. For these experiments, we switched to an even smaller, more idealized problem. Slowly-Changing Regression is a computationally inexpensive problem where we can run a single experiment on a CPU core in 15 minutes, allowing us to do extensive studies. As the name suggests, this is a regression problem—meaning the labels are real numbers, with a square loss, rather than nominal values with a cross-entropy loss—and the nonstationarity is slow and continual rather than abrupt as in a switch from one task to another. In Slowly-Changing Regression, we study loss of plasticity for networks with six different activation functions: sigmoid, tanh, ELU (Clevert et al., 2016), leaky-ReLU (Maas et al., 2013), ReLU (Nair & Hinton, 2010), and Swish (Ramachandran et al., 2018).
Figure 4: The target function and the input in the Slowly-Changing Regression problem. The input has $m+1$ bits. One of the flipping bits is chosen after every $T$ time steps, and its value is flipped. The next $m-f$ bits are i.i.d. at every time step. And, the last bit is always one. The target function is represented by a neural network with a single hidden layer of LTUs. Each weight in the target network is $-1$ or $1$.
In Slowly-Changing Regression, the learner receives a sequence of examples. The input for each example is a binary vector of size $m+1$. The first $f$ bits in the input vector change slowly. After every $T$ example, one of the first $f$ bits is chosen randomly, and its value is flipped. These first $k$ bits remain fixed for the next $T$ examples. The parameter $T$ allows us to control the speed at which the input distribution changes. The next $m-f$ bits are randomly sampled for each example. Lastly, the $(m+1)$-th bit is a bias term with a constant value 1.
The target output is generated by running input through a neural network, which is set at the start of the experiment and kept fixed. As this network generates the target output and represents the desired solution, we call it the target network. The weights of the target networks are randomly chosen to be +1 or -1. The target network has one hidden layer with a Linear Threshold Unit (McCulloch & Pitts, 1943), or LTU, activation. The value of the $i$-th LTU is one if the input is above a threshold $\theta_{i}$, and zero otherwise. $\theta_{i}$ is set equal to $(m+1)\cdot\beta-S_{i}$, where $\beta\in[0,1]$ and $S_{i}$ is the number of input weights with negative value (Whitehead and Sutton, 1994). The details of the input and target function in the slowly-changing regression problem are also described in Figure 4.
The details of the specific instance of the slowly-changing regression problem we use in this paper and the learning network used to predict its output are listed in Table 4. Note that the target network is more complex than the learning network as the target network is wider, with 100 hidden units, while the learner has just five hidden units. Thus, because the input distribution changes every $T$ example and the target function is more complex than what the learner can represent, there is a need to track the best approximation.
Table 2: Implementation details for the slowly-changing regression problem and the learning network
slowly-changing regression Problem Parameters
Parameter Name Description Value
$m$ Number of input bits 21
[…]
Learning Network Parameters
Parameter Name Value
Number of hidden layers 1
[…]
We applied learning networks with different activation functions to the slowly-changing regression. The learner used the backpropagation algorithm to learn the weights of the network. We used uniform Kaiming distribution (He et al., 2015) to initialize the learning network’s weights. The distribution is $U(-b,b)$ with bound, $b=gain*\sqrt{\frac{3}{num\_inputs}}$, where $gain$ is chosen such that the magnitude of inputs does not change across layers. For tanh, Sigmoid, ReLU, and Leaky-ReLU, $gain$ is 5/3, 1, $\sqrt{2}$, and $\sqrt{2/(1+\alpha^{2})}$ respectively. For ELU and Swish, we used $gain=\sqrt{2}$, as was done in the original works by Clevert et al. (2016) and Ramachandran et al. (2018).
We ran the experiment on the Slowly-changing regression problem for 3M examples. For each activation and value of step size, we performed 100 independent runs. First, we generated 100 sequences of examples (input-output pairs) for the 100 runs. Then these 100 sequences of examples were used for experiments with all activations and values of the step size parameter. We used the same sequence of examples to control the randomness in the data stream across activations and step sizes.
The results of the experiments are shown in Figure 5. We measured the squared error, the square of the difference between the true target and the prediction made by the learning network. In Figure 5, the squared error is presented in bins of 40k examples. This means that the first data point is the average squared error on the first 40k examples, the next is the average squared error on the next 40k examples, and so on. The shaded region in the Figure shows the standard error of the binned error.
Figure 5: Loss of plasticity is robust across different activations.
Figure 5 shows that in Slowly-changing regression, after performing well initially, the error increases for all step sizes and activations. For some activations like ReLU and tanh, loss of plasticity was severe, and the error increased to the level of the linear baseline. While for other activations like ELU, loss of plasticity was less severe, but still there was a significant increase in the error. These results mean that the loss of plasticity is not resolved by using different activations.
So far in the paper, we performed direct tests of loss of plasticity with backpropagation. Our experiments in Continual Imagenet, online Permuted MNIST, and Slowly changing regression show that loss of plasticity is a general phenomenon, and it can be catastrophic in some cases. It happens for a wide range of hyper-parameters, activation functions, rates of distribution change, and for both under and over-parameterized networks. Although continual learning has been studied for a long time from (Caruana, 1997; Ring, 1998; Parisi et al. (2019)), a direct test of the ability to maintain plasticity was missing. The experiments so far in the paper fill this gap and show that conventional backpropagation does not work in continual learning.
There are two major goals in continual learning: maintaining stability and maintaining plasticity. Maintaining stability is concerned with memorizing useful information, and maintaining plasticity is about finding new useful information when the data distribution changes. It is difficult for current deep-learning methods to remember useful information as they tend to forget previously learned information (McCloskey & Cohen, 1989; French, 1999; Parisi et al., 2019). Various ingenious methods have been proposed to tackle the forgetting issue (Kirkpatrick et al., 2017; Aljundi et al., 2019; Riemer et al., 2019; Javed & White, 2019; Veniat et al., 2021), but it remains an unsolved problem. We focused on continually finding useful information, not on remembering useful information. Our work on loss of plasticity is different but complementary to the work on maintaining stability.
Table 2: Implementation details for the slowly-changing regression problem and the learning network